a pipe 100 feet long .75id what is the total amount of liquid in us galons will it hold
Answer by Alan3354(69443)

a pipe 100 feet long .75id what is the total amount of liquid in us galons will it hold
----------
L = 1200 inches
r = 0.375 inches = 3/8"
------------
Vol = pi*r^2*L
V = pi*(9/64)*1200 = 10800pi/64 = 168.75pi cubic inches
V = 168.75pi/231 gallons
V =~2.295 gallons